OpenStudy (jacksonjrb):

Given one or more of its roots, complete the solution set for each equation.

OpenStudy (jacksonjrb):

\[x^4-\frac{ 13 }{ 2 }x^3+8x^2-\frac{ 5 }{ 2 }x=0\]